OpenStudy (anonymous):

use radical expression to write the expression: 7m^1/4

OpenStudy (anonymous):

\[\sqrt[4]{7m}\]

OpenStudy (anonymous):

1/4 = 4th root \[7\sqrt[4]{m}\]

OpenStudy (anonymous):

Depends on your parentheses.

OpenStudy (saifoo.khan):

\[\sqrt[4]{7m}\]